OpenStudy (crashonce):

Suppose the number of a kind of bacteria on the nth day (A) can be calculated by the following formula: A=k log n, where k is a constant. If there are 2000 bacteria on the 100th day, when will the amount attain 4000?

OpenStudy (crashonce):

@ganeshie8

OpenStudy (crashonce):

@ikram002p

OpenStudy (ikram002p):

soo A(100)=k log(100)=2000 thus find k :D

OpenStudy (anonymous):

Boltzmann ENtity: S = k log V

OpenStudy (ikram002p):

k=1000/log 10 then A(n)=k log n =4000 solve for n

OpenStudy (crashonce):

so its 1000

OpenStudy (crashonce):

@ikram002p so the answer is 10000th day?

OpenStudy (ikram002p):

yeah k=1000 next step solve for n A(n)=1000 log n =4000

OpenStudy (crashonce):

so is it 10000

OpenStudy (ikram002p):

yes :) n=10^4

OpenStudy (anonymous):

A = n^k 2000= (100)^k \[k \approx 1.6505\]

OpenStudy (crashonce):

thanks @ikram002p

OpenStudy (ikram002p):

np :)

OpenStudy (anonymous):

4000 = x^1.6505 x = 152 days

OpenStudy (ikram002p):

how is that adjax ?

OpenStudy (anonymous):

I just converted the expression into exponential form from the log one: A = n^k 2000 = (100)^k k = 1.6505 when A = 4000 n = ? k =1.6505{constant} 4000 = n^1.6505 log 4000 = 1.6506 log n ..... n = 152 days(rounded-off) @ikram002p

OpenStudy (ikram002p):

A=k log n A=log (n^k) 10^A=n^k

OpenStudy (anonymous):

a blunder on my part :k = 2.1505 a little modification and the answer will come currect: Answer is : n= 47 days